### v1

# fMRI-3D: A Comprehensive Dataset for Enhancing fMRI-based 3D Reconstruction

Abstract
Reconstructing 3D visuals from functional Magnetic Resonance Imaging (fMRI) data, introduced as Recon3DMind in our conference work [1], is of significant interest to both cognitive neuroscience and computer vision. To advance this task, we present the fMRI-3D dataset, which includes data from 15 participants and showcases a total of 4,768 3D objects. The dataset comprises two components: fMRI-Shape, previously introduced and accessible at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Shape, and fMRI-Objaverse, proposed in this paper and available at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse. fMRI-Objaverse includes data from 5 subjects, 4 of whom are also part of the Core set in fMRI-Shape, with each subject viewing 3,142 3D objects across 117 categories, all accompanied by text captions. This significantly enhances the diversity and potential applications of the dataset. Additionally, we propose MinD-3D, a novel framework designed to decode 3D visual information from fMRI signals. The framework first extracts and aggregates features from fMRI data using a neuro-fusion encoder, then employs a feature-bridge diffusion model to generate visual features, and finally reconstructs the 3D object using a generative transformer decoder. We establish new benchmarks by designing metrics at both semantic and structural levels to evaluate model performance. Furthermore, we assess our model’s effectiveness in an Out-of-Distribution setting and analyze the attribution of the extracted features and the visual ROIs in fMRI signals. Our experiments demonstrate that MinD-3D not only reconstructs 3D objects with high semantic and spatial accuracy but also deepens our understanding of how human brain processes 3D visual information. Project page at: https://jianxgao.github.io/MinD-3D.
Index Terms:
FMRI decoding, 3D vision, Dataset, Diffusion model.1 Introduction
Functional Magnetic Resonance Imaging (fMRI), a kind of signal that can be obtained in a non-invasive way, could capture blood changes in the human brain induced by neuronal activity. Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities. Some recent studies [2, 3, 4, 5, 6] have successfully reconstructed high-quality images from fMRI signals by utilizing powerful generative models [7, 8]. These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
Existing methods primarily focus on reconstructing 2D visual information, but the human visual system extends far beyond merely processing flat images. It possesses the extraordinary ability to transform 2D projections into rich 3D representations. This complex mechanism allows us to perceive the world in depth, recognizing attributes like size, distance, and spatial depth. In contrast to previous studies, our research centers on modeling the brain’s 3D visual capabilities. We introduce a new task, called Recon3DMind (Reconstructing 3D Objects from Mind), which leverages advanced computer vision techniques to decode and reconstruct the 3D visual information perceived by the brain from fMRI signals. This task goes beyond merely extracting semantic features, incorporating spatial and structural dimensions that are essential for a comprehensive understanding of 3D vision.
Several studies [9, 10, 11] have demonstrated that the brain’s mechanisms for 3D visual perception are significantly more intricate than those for 2D perception. This complexity is reflected in the distinct activation of brain regions during 3D visualization tasks [12, 13]. As a result, relying solely on semantic features is insufficient to fully model the brain’s capacity for 3D spatial perception. Effectively describing 3D objects requires taking into account not only their semantic features but also their shape and structural properties. For instance, two cars may appear identical when viewed head-on, yet differ greatly in length when viewed from the side. This example underscores the importance of capturing the full range of spatial and structural features to authentically represent 3D objects. Accordingly, our work seeks to advance the modeling of human 3D perception by developing an enhanced fMRI feature extractor. This extractor is designed to capture semantic elements, spatial structures and other 3D-specific characteristics from fMRI signals. This approach aims to enable a more complete and accurate reconstruction of 3D visual information.
In our conference work [1], we introduced the fMRI-Shape dataset to tackle the significant challenge of the lack of datasets pairing fMRI data with 3D visuals for this complex task. The dataset comprises data from 14 participants and 1,624 3D objects. However, the Core set was limited to just 13 categories of 3D objects. To address this limitation in category diversity and to expand the number of objects, we propose fMRI-Objaverse, which includes data from 5 participants and 3,142 3D objects across 117 categories, accompanied by text captions. Notably, it shares 4 participants with the Core set of fMRI-Shape, significantly enhancing the diversity of fMRI-Shape, as shown in Fig. 2. We collectively refer to these two datasets as fMRI-3D, aiming to support various experimental setups and further promote research within the community.
During fMRI data collection, we present 3D objects through 360-degree view videos, providing comprehensive visualizations that stimulate the brain’s perception of 3D objects and facilitate the collection of high-quality data. In our approach, participants watch 360-degree videos of stationary 3D objects from ShapeNet [14] and Objaverse [15], where a rotating camera completes a full orbit around each object, offering a complete view from all angles. This method ensures detailed and accurate capture of fMRI signals, as participants engage with the objects, allowing for the full range of spatial features to be recorded. As shown in Fig.3, we also analyze the variation in fMRI data across both subjects and objects. Interestingly, the variation across subjects is even greater than that across objects. After careful preprocessing, these recordings are transformed into multi-frame fMRI signals, resulting in a rich dataset for detailed analysis. The complexities and specific features of the fMRI-3D dataset will be discussed further in Sec. 3.
Leveraging our carefully curated fMRI-3D dataset, we introduce an innovative and efficient three-stage framework called MinD-3D to extract both spatial structure and semantic features from multi-frame fMRI signals and reconstruct the corresponding 3D visual stimuli.
In the first stage, we utilize an encoder [6] pre-trained on NSD [16] dataset to extract spatial features from the fMRI data. These features are then aggregated across multiple frames using a feature aggregation module. To ensure the biological relevance and effectiveness of these extracted features, we align them with the visual-spatial characteristics of the corresponding objects using contrastive learning loss. This alignment ensures that the extracted features are both accurate and biologically meaningful in relation to the brain’s visual processing mechanisms.
The second stage focuses on decoding the brain’s visual activity from the fMRI data. We achieve this through a transformer-based diffusion model trained in feature space, conditioned on the fMRI-derived features. This conditioning enables us to progressively generate accurate visual representations, translating complex brain activity into understandable visual data.
In the final stage, we focus on reconstructing the 3D models as they are perceived by the human brain. To do this, we train a latent-adapted Argus [17] model, which uses the visual features generated in the second stage along with the initial fMRI features to create 3D objects. This process mimics how the brain’s visual cortex operates, providing a simulated insight into the brain’s perception and processing of 3D objects.
To measure the effectiveness of our model, we design new benchmarks that evaluate its performance at both the semantic and structural levels. These metrics comprehensively assess the model’s ability to generate 3D representations that are structurally and semantically accurate. We evaluate our model under both standard and Out-of-Distribution settings, and it consistently outperforms baseline models. Additionally, we conduct in-depth analyses of our fMRI-3D dataset and the features extracted by MinD-3D, exploring how the brain perceives different angles, objects, and semantic information within specific ROIs (regions of interest). We further validate the relevance of our model’s features by correlating them with brain regions, confirming that the representations it produces are consistent with the brain’s visual information processing.
This paper expands upon our preliminary conference work, and we summarize the key contributions as follows:
We introduce fMRI-Objaverse, a large-scale extension of the original dataset. Together with fMRI-Shape, we collectively refer to these datasets as fMRI-3D, designed to support various experimental setups and advance research in the field.
We establish new benchmarks for the task of 3D visual reconstruction from human brain data.
We conduct extensive experiments to analyze the contributions of our proposed dataset in decoding fMRI signals, further validating the effectiveness of MinD-3D.
2 Related work
2.1 fMRI Decoding Methods
Current fMRI decoding methods primarily focus on reconstructing the vision perception in a 2D format, such as the images or videos perceived by humans. This is a challenging task, as it involves extracting relevant features from fMRI signals with precision to recreate accurate 2D representations. Deep learning methods, known for their impressive capabilities, are particularly suited to address this challenge. Initial successes in this area have been demonstrated by earlier methods [18, 19, 20]. Subsequent studies [21, 22] have shown that generative models are particularly effective for these tasks, leading to the employment of various diffusion models [2, 3, 4, 23] as decoders to reconstruct visual scenes, achieving remarkable results. However, these studies have been limited to 2D visual representations and the related vision ROIs. In this paper, we aim to extend the scope of fMRI visual decoding to 3D representations, involving more vision ROIs. Our goal is to directly reconstruct 3D objects from fMRI signals. To accomplish this, we propose a new framework that employs a transformer-based feature encoder for extraction and aggregation. This framework translates neural space data into visual space and utilizes a powerful 3D decoder to reconstruct the 3D object, leveraging features from the visual space.
2.2 Diffusion Models
Diffusion models [24, 25] are exceptional generative tools for both pixel and feature generation. As a variant, the latent diffusion model [7], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data itself. Dit [26] replaces the backbone of diffusion models with transformers, which will improve the performance and scalability of these models. This approach, operating in the latent space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space. In this paper, we aim to leverage the potent feature-generation capabilities of diffusion models to generate visual features based on fMRI features. To achieve this, we adapt a transformer-based diffusion model, focusing solely on its latent component. The conditional information driving the model is derived from the fMRI features.
2.3 3D Generation
3D generation can be accomplished through various methods [17, 27, 28]. Some methods [29, 27] employ 3D Gaussian splatting [30] for this purpose. Other studies [31, 32] utilize diffusion models to generate multi-view representations of objects, subsequently constructing 3D models. Additionally, traditional and direct approaches leverage autoregressive methods [33, 34, 17] for 3D object generation. In our study, we adapt Argus [17], a robust 3D generative model with several transformer layers, as our decoder to generate 3D objects from fMRI data. This approach integrates visual features generated by the preceding diffusion module. These visual features serve as conditional embeddings for Argus. This synergistic integration aims to enhance the model’s ability to accurately reconstruct 3D objects from complex brain activity.
3 Experimental Designs and Curated Dataset
In this section, we detail the procedures for collecting the proposed fMRI-3D dataset, which consists of two components: fMRI-Shape and fMRI-Objaverse. The scale of fMRI-3D is compared to other benchmark datasets, including NSD [16], BOLD5000 [35], GOD [18], and Video-fMRI [19], as illustrated in Fig. 4. Specific details about fMRI-Shape and fMRI-Objaverse are provided in Tab. I. For all experiments, written informed consent was obtained from each participant, and the study was approved by the ethical review board. To better illustrate brain activation patterns and demonstrate the utility of the fMRI-3D dataset, we analyze and visualize responses to three distinct objects across six subjects, as shown in Fig. 3. Note that only voxels with activation levels above the 50th percentile are displayed. We also compute the variation across subjects and objects, with red and blue regions indicating higher activation values, respectively, reflecting areas in the human brain sensitive to the stimuli. This visualization highlights significant individual differences in brain activation across subjects, which are more pronounced than the variations in responses to different objects. These findings emphasize the inherent challenges and underscore the importance of the AP and APAC settings. All participants had normal or corrected-to-normal vision. The fMRI-3D dataset will be made publicly available to support further research in Recon3DMind.
| Participant | Males/Females | Category | Objects | Frames | |
| fMRI-Shape | 14 | 7/7 | 55 | 1624 | 123200 |
| Core Set | 8 | 4/4 | 13 | 1404 | 14040 |
| AP Set | 2 | 1/1 | 13 | 104 | 1040 |
| APAC Set | 4 | 2/2 | 55 | 220 | 2200 |
| fMRI-Objaverse | 5 | 2/3 | 117 | 3142 | 125680 |
| fMRI-3D (Total) | 15 | 7/8 | 172 | 4768 | 248880 |
3.1 fMRI-Shape
FMRI-Shape contains data from 14 participants who were unaware of the objectives of the work. To ensure diversity in the dataset, the 3D objects were sourced from ShapeNetCore [14], which includes 55 object categories. We employed the rendering technique from Zero123 [36] to render 192 images using Blender and generated 8-second videos at 24 fps for each object. These videos depict the 3D objects rotating 360 degrees at a 60-degree pitch angle, as illustrated in Fig. 6. The dataset is available for download at: https://huggingface.co/datasets/Fudan-fMRI/fMRI-Shape.
1) Core Set: The core set of fMRI-Shape includes data from 8 participants (4 males and 4 females, aged 21 to 29, Participants No. 1-8). A total of 1,404 objects were selected from 13 commonly used categories in 3D reconstruction literature [37, 38, 39] within ShapeNetCore. For each category, 100 objects were used for training and 8 for testing, resulting in 108 objects per category. For more details, please refer to our conference version [1].
2) Across-Person Set (AP Set): The AP Set was designed for Out-of-Distribution (OOD) testing and includes fMRI data from 2 participants (1 male aged 24 and 1 female aged 26, Participants No. 9 and 10) who viewed the test objects from the Core set.
3) Across-Person & Across-Class Set (APAC Set): The APAC Set presents a more challenging OOD test compared to the AP Set. It includes data from 4 participants (2 males and 2 females, aged 22 to 26, Participants No. 11-14). For this set, we randomly selected 4 objects from each of the 55 categories in ShapeNetCore, distinct from those in the Core set, resulting in a total of 220 objects.
As illustrated in the middle part of Fig. 6, individual differences among participants pose significant challenges for generalization. The AP and APAC sets are crucial for OOD testing and will serve as important benchmarks for assessing the generalization capability of 3D decoding models.
3.2 fMRI-Objaverse
FMRI-Objaverse, as partially shown in Fig. 5, includes data from five participants, all of whom were unfamiliar with the study. Four participants (2 males and 2 females, aged 22 to 26, identified as Nos. 1, 6, 7, and 8) overlap with the core fMRI-Shape dataset, providing an important extension in terms of diversity and scale. Additionally, the fifth participant (No. 15), a 22-year-old female, was included to further expand the dataset. To enhance our dataset, we selected 3,142 objects from the top 117 object categories in Objaverse [15], based on a subset filtered by LGM [27] and enriched with text descriptions from Cap3D [40]. Unlike in fMRI-Shape, each 3D object in this dataset was rendered into 384 frames, generating a 6.4-second video at 48 fps using Blender. Each participant spent approximately 8 hours in experimental sessions, divided into 53 sessions. During each session, participants viewed 60 videos in a randomized order, except for the last session, with 1.6-second rest intervals between each pair of objects. To prevent low-quality data due to visual fatigue, we randomly reversed the rotation direction for 40% of the selected objects. All objects were presented once to each participant. (Note: Our MRI machine samples data every 800ms, so we selected 6.4-second videos with a 1.6-second rest period between them.) This extension supports further multimodal experiments and applications. The objects in Objaverse contain more detail and a wider variety of categories, and the higher fps videos present stronger visual effects, posing a significant challenge for reconstructing them in fMRI-Objaverse. The dataset is available for download at: https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse.
3.3 Data Acquisition and Preprocessing
T1 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle ). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle , multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 1.25Hz, so each video segment corresponds to a total of 10 frames of task-state fMRI signals.
Stimuli were presented using an LCD screen () positioned at the head of the scanner bed. Participants viewed the monitor via a mirror mounted on the RF coil and fixated a red central dot (0.4° × 0.4°).
Preprocessing was performed using fMRIPrep [41, 42]. Following [5], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Given the delay of the BOLD signal by 6 seconds, we applied z-scoring to the data points across every vertex within each run, incorporating a 6.4-second lag. These normalized values were then projected onto 1023 × 2514 pixel 2D images using pycortex. For analysis, Regions of Interest (ROIs) were selected from the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) atlas in the 32k_fs_LR space. These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”. Subsequently, the ROIs were converted into a 256 × 256 image, as illustrated in the right part of Fig. 6.
4 Problem Setup
Recon3DMind tackles a critical challenge in cognitive neuroscience: developing computational models that can accurately interpret and reconstruct the brain’s 3D visual comprehension. This endeavor not only bridges the gap between cognitive neuroscience and computer vision but also has the potential to advance the latter field in unprecedented ways. In this paper, we focus on the specifics of fMRI-based 3D reconstruction, providing detailed definitions and formulas that underpin our approach.
We begin with the acquisition of a multi-frame fMRI signal, denoted as , where (with or ). These signals correspond to both a 3D object mesh, , and a video, , with (where or frames), which the subject observes. The task requires an efficient encoder, , capable of extracting both spatial structural and semantic features from the fMRI signal. It is important to note that while a single frame of fMRI data is sufficient to extract semantic information for 2D image reconstruction, reconstructing 3D structures requires additional spatial structural features. Therefore, multiple frames of fMRI data are input to capture these comprehensive spatial features from the spatio-temporal signals. Mathematically, this is expressed as: .
After feature extraction, a powerful decoder is used to reconstruct the original 3D mesh, , based on the extracted feature : . Thus, our model can be succinctly described as , where the transformation is represented as .
To effectively implement this model, it is crucial to leverage both the fMRI signals and the corresponding video to train the model . For this, we propose a three-stage, innovative, and efficient framework. Each stage is carefully designed to capture different aspects of the fMRI data and the associated visual stimuli, ensuring a comprehensive and accurate 3D reconstruction from the complex neural signals. This process not only pushes the boundaries of current computer vision techniques but also provides valuable insights into how the human brain processes 3D spatial information.
5 Method
5.1 Overview
As depicted in Fig. 7, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames. 2) Feature Bridge Diffusion Model: This module generates visual features based on the features extracted by the Encoder. 3) Latent Adapted Decoder: Finally, this component recovers 3D objects, using both the fMRI features and the generated visual features.
5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [43, 44, 6], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [45]. Our encoder, termed the Neuro-Fusion Encoder (NFE), employs the encoder of LEA [6] as , along with a Feature Aggregation module (), to transform fMRI frames into embeddings. Inspired by contrastive learning [46, 47], we then align these embeddings with the vision space through CLIP [46].
Initially, we process each fMRI frame in parallel to obtain the spatial fMRI embeddings:
Then, we use the Feature Aggregation module to aggregate the fMRI embeddings into fMRI latent feature :
To align the fMRI latent feature with the vision space, we employ the ViT-B/32 CLIP vision encoder and here to encode video frames into vision embeddings. During training, we randomly select n frames as and compute the average vision embeddings as the visual latent feature :
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between fMRI features and visual features :
During training, we optimize the neuro-fusion encoder initialized with pre-trained weights, while the CLIP vision encoder remains frozen. It’s worth noting that the CLIP vision encoder and images are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features and embedding aligned with the visual space. The fMRI features will be used as the conditional information for the feature bridge diffusion model.
5.3 Feature Bridge Diffusion Model
This section describes the Neural-Visual Synthesis process via a transformer-based diffusion model, herein referred to as the Feature Bridge Diffusion Model (FBDM), inspired by the hierarchical design proposed in [48]. The FBDM’s primary role is to bridge fMRI latent features with their visual counterparts , effectively transforming neurological signals into visual representations. In the forward diffusion process, we treat the visual latent feature as and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [4]:
During the reverse diffusion process, FBDM is tasked with predicting and reversing the introduced noise to reconstruct the latent visual feature. This is achieved by estimating the noise at each timestep, conditioned on the fMRI latent feature :
Differing from U-Net based models, our approach utilizes a transformer to leverage the self-attention mechanism for integrating the conditional information . This is done by embedding the timestep , the noise-perturbed feature , and the conditional fMRI feature into a unified representation, further augmented with a learnable token to capture the target vision features accurately. The objective of FBDM training is to minimize the discrepancy between the actual noise and its prediction:
Upon training completion, the model discards the initial visual latent feature and enters a generative phase. It begins with pure noise and progressively refines this through the reverse diffusion process to generate the predicted visual latent feature . This process effectively translates the aligned fMRI latent features into visual representations, serving as essential conditional information for the subsequent latent-adapted decoding phase.
5.4 Latent Adapted Decoder
The Latent Adapted Decoder (LAD) leverages Argus [17], a state-of-the-art 3D mesh generation method, as its backbone. This component is tasked with generating the Vector Quantized (VQ) latent representation of target 3D objects from conditional information processed through a generative auto-regressive transformer. The resulting 3D objects are then reconstructed using the VQ decoder.
To tailor Argus for processing fMRI input and to retain its generative efficiency, we employ adapter tuning during training. This involves integrating the adapter layers within the transformer block. Following classic methods [49, 50], we apply two feed-forward linear layers to effectively adapt the self-attention outputs. The adapter layer is systematically inserted at every transformer blocks, ensuring that the model remains adaptable and efficient. The decoder formulates the VQ latent vector for the target object as follows:
Subsequently, the 3D VQ-decoder decodes this latent vector into the final 3D mesh representation :
The decoder aims to minimize the negative log-likelihood loss during optimization:
For training, the decoder is fine-tuned from pre-trained weights, with the trained neuro-fusion encoder and the feature bridge diffusion model. This strategy preserves Argus’s generative capabilities while adapting to new inputs and minimizing the computational overhead associated with training.
| Methods | Semantic-Level | Structure-Level | |||||
| 2-way | 10-way | LPIPS | SSIM | FPD | CD | EMD | |
| LEA-3D [6] | 0.787 | 0.371 | 0.527 | 0.562 | 4.229 | 2.291 | 5.347 |
| fMRI-PTE-3D [5] | 0.815 | 0.392 | 0.433 | 0.694 | 3.571 | 1.992 | 4.621 |
| w/o Both | 0.789 | 0.367 | 0.479 | 0.616 | 3.694 | 2.205 | 5.073 |
| w/o Diffusion | 0.801 | 0.385 | 0.423 | 0.669 | 3.526 | 2.071 | 4.625 |
| w/o Contrastive | 0.823 | 0.419 | 0.319 | 0.701 | 3.315 | 1.826 | 4.027 |
| MinD-3D (full) | 0.839 | 0.432 | 0.230 | 0.734 | 3.157 | 1.742 | 3.833 |
6 Experiments and Benchmark
To establish new benchmarks for 3D visual decoding, we conduct experiments in both standard and Out-of-Distribution (OOD) settings. In this section, we introduce the metrics and provide details of the experiments.
6.1 Metrics
To effectively evaluate the performance of our models on this novel task, we employ metrics across two key dimensions: semantic-level and structure-level.
Semantic Level. To assess the semantic quality of our model, we adopt standard metrics commonly used in prior 2D fMRI studies [2, 3, 51, 52, 23], specifically N-way top-K accuracy. We report 2-way-top-1 and 10-way-top-1 accuracy, as shown in Table II. Additionally, we calculate the Learned Perceptual Image Patch Similarity (LPIPS) [53] to evaluate the perceptual quality of the semantic information. These metrics are computed by comparing the reconstructed images with the ground truth (GT) images. Both the reconstructed and GT objects are rendered into images at every 60-degree rotation, and the metrics are calculated for each frame. The final score is derived by averaging the values across frames.
Structure Level. In addition to semantic evaluation, it is essential to measure how well our model captures the geometrical structure of objects. We utilize common 3D reconstruction metrics [54, 55, 17]: Fréchet Point Cloud Distance (FPD, scaled by ), Chamfer Distance (CD, scaled by ), and Earth Mover’s Distance (EMD, scaled by ). These metrics are computed by sampling point clouds from the GT and reconstructed meshes. Additionally, we calculate the frame-averaged Structural Similarity Index (SSIM) using the same rendering process to evaluate similarity from various viewpoints.
6.2 Implementation Details
As detailed in Sec. 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and 192 corresponding images. For the experiment for fMRI-Shape, to maximize dataset usage and incorporate data augmentation, we randomly select 6 fMRI frames for training. The original 1023 × 2514 2D fMRI images are resized to 256 × 256 for processing. For the image data, 8 frames are randomly selected for training, with each image resized to 224 × 224. During inference, only the middle 6 fMRI frames are used, without additional images. Regarding the model architecture, we set to incorporate adapter layers.
The model training process is divided into two stages. In the first stage, we optimize the neuro-fusion encoder and the feature bridge diffusion model. The neuro-fusion encoder is initialized with pre-trained weights, while the feature bridge diffusion model is trained from scratch, which takes approximately 2.5 days on a single A100 GPU. In the second stage, all parameters, including those of the adapters, are optimized, while the parameters in Argus are frozen. This stage is completed in roughly one day on a single A100 GPU.
6.3 Standard Experiment
As the first attempt to model 3D imaging within the human brain, we define the standard experiment setting as training and testing on the pre-split, person-specific Core set in fMRI-Shape, using the defined metrics to evaluate our model’s performance. Given the complexity of this task, which involves various brain regions, direct comparisons with existing models are not feasible. To reduce training costs, we adapted the LEA and fMRI-PTE models, which were trained on the same vision ROIs and have strong performance. These models utilize the same 3D decoder as MinD-3D. These adaptations serve as baselines in this study, enabling a more contextual and fair comparison within this novel domain.
Tab. II reports the averaged metrics at both the structural and semantic levels for all subjects. MinD-3D demonstrates superior performance across all metrics, achieving approximately 83.9% accuracy in 2-way-top-1, 43.2% accuracy in 10-way-top-1, and scores of 0.734, 0.230, 3.157, 1.742, and 3.833 in other metrics. These results indicate that our model not only excels in generating objects with high semantic accuracy but also performs exceptionally well in preserving structural similarity, outperforming baseline methods. Tab. II also highlights the effectiveness of both the diffusion model and the contrastive learning module, which play critical roles in the success of MinD-3D.
Qualitative results from LEA-3D, fMRI-PTE-3D, and our method are presented in Fig. 8. MinD-3D consistently produces 3D objects that are structurally similar to their real counterparts, while maintaining semantic integrity in most cases. This highlights the robustness of our model in handling this challenging task and its ability to generate faithful reconstructions.
Furthermore, given the individual differences highlighted in Fig. 3, we also examine our model’s performance on each subject to assess its generalization ability. Fig. 9 shows several examples from 5 different subjects, demonstrating the effectiveness of the MinD-3D model. Due to varying brain responses among subjects to the same objects, some differences are observed in the results. However, overall, the model successfully reconstructs both the category and characteristics of the objects. This validates the feasibility of the proposed task, Recon3DMind, and the associated fMRI-Shape dataset developed for this study.
6.4 Out-Of-Distribution Experiments
To effectively utilize a subset of the fMRI-Shape dataset and further assess the generalization capabilities of our proposed MinD-3D model, we conduct two Out-Of-Distribution (OOD) experiments under challenging settings: 1) Across-Person Testing (APT): In APT, we evaluate our model, which was trained only on Subject 1, using the data from Subject 9. We compare the results with the baselines and report the metrics in Tab. III. 2) Across-Person & Across-Class Testing (APACT): In APACT, we similarly evaluate our model, trained solely on Subject 1, with the data from Subject 11. We also compare with the baselines and report the metrics in Tab. III.
We present the reconstructed objects from AP & APAC testing in Fig. 10. As shown in Fig. 3, individual differences significantly impact the results, which our AP & APAC tests empirically confirm. Despite the high difficulty, MinD-3D successfully recovers the basic shapes of the objects, providing a strong baseline for the community. While performance in these OOD scenarios does not reach In-Distribution (ID) levels—an expected outcome given the task’s complexity and the substantial individual differences and domain gaps—our method still surpasses existing baselines. This demonstrates the robustness of MinD-3D and establishes a new benchmark for future work.
| Methods | APT | APACT | ||||
| FPD | CD | EMD | FPD | CD | EMD | |
| LEA-3D | 5.362 | 3.627 | 6.174 | 6.958 | 4.944 | 8.107 |
| fMRI-PTE-3D | 4.501 | 2.956 | 5.772 | 6.261 | 4.570 | 7.843 |
| MinD-3D | 3.838 | 2.415 | 5.117 | 5.689 | 4.181 | 7.194 |
7 Analysis
7.1 Voxel Importance Analysis for Object Angle Variations
In this section, we explore the pattern when the angle of the object changes. Most directly, we perform linear regression on the whole-brain fMRI signal data for classification. The weight of each voxel in the brain serves as its importance score. Specifically, we performed logistic regression on the whole-brain fMRI signals using all voxels from subject 1 while viewing objects from two different angles (the 3rd and 9th frames of fMRI). We randomly selected 70% of the data for training and used the remaining 30% for testing. The accuracy of the linear classifier is 91.34%. To evaluate the importance of each voxel, we visualized the absolute values of the classifier weights in Fig. 11. Red regions indicate voxels with higher coefficients, while others represent lower coefficients. These ROIs include “2, AIP, FFC, FST, IPS1, LIPv, LO3, MIP, MST, MT, PH, PFt, PGp, PHT, TPOJ2, V1, V2, V3, V3A, V3B, V3CD, V4, V4t, V7, V8, VIP, 7PC, and 7PL.” These ROIs involve the parietal and occipital lobe, corresponding to spatial information processing and visual processing in the human brain, respectively.
7.2 Analysis of Different Objects
In this section, we explore the brain activity patterns associated with viewing different objects. The analysis is divided into two parts:
7.2.1 Between Different Types of Objects
Building on the previous angle-based experiment, we perform logistic regression on the whole-brain fMRI signals of subject 1 to differentiate between cars and rifles. The linear classifier achieved an accuracy of 79.54%. We also visualize the absolute values of the classifier weights in Fig. 12, where red regions indicate voxels with higher coefficients and other colors represent lower coefficients. The regions of interest (ROIs) include “25, 47s, A5, AIP, PreS, STGa, STSda, TGd, V1, V2, V3, V3A, V4, V4t, V6, V7, and V8,” primarily located in the temporal and occipital lobes, corresponding to high-level visual functions and visual processing, respectively.
7.2.2 Between Objects Within the Same Category
In this case, classification experiments are more challenging due to the lack of specific labels for supervision. However, we visualize the absolute difference between two objects (a car and a plane) in Fig. 13 for analysis. In the figure, deep blue indicates voxels with higher values, while lighter colors represent lower values. Red rectangles highlight regions with higher values. Notably, these regions are predominantly located in the parietal and occipital lobes, which are known to play key roles in object differentiation.
7.3 Explore how our brain understand semantic information
Our proposed MinD-3D successfully reconstructs 3D objects both semantically and structurally, so we aim to explore how the brain processes semantic information of different objects. Our model mainly focuses on the visual regions and does not include a specific branch for semantic information. To achieve this, we use Class Activation Mapping (CAM) methods [56] to analyze the importance of each part of the input fMRI frame. We We present CAMs for three objects in Fig. 14. From this, we identify the following ROIs that may be involved in processing visual semantic information: FFC, FST, IP0, MT, MST, PHA1, PHA2, PHA3, PH, PIT, V1, V2, V3, V3A, V4, V4t, V6A, V7, V8, VMV1, VMV2, VMV3, and VVC.
7.4 Features Analysis
To understand the biological relevance of our feature extractor, in line with previous studies [19, 57, 58], we employ a linear encoding model to project the extracted fMRI features onto brain activity patterns, using subject 1 as a representative case for analysis. Specifically, we utilize ridge regression to establish the mapping between and the fMRI signals within the training dataset. We then project from the test set into the fMRI space and compute the Pearson correlation coefficient to assess the relationship. As illustrated in Fig. 15, areas positively correlated with visual Regions of Interest (ROIs) are marked in red, while other regions showing positive correlations are highlighted in blue. The results demonstrate a strong positive correlation between and visual ROIs, confirming the effective extraction of relevant features. Additionally, we perform a similar correlation analysis with the visual latent feature , which also shows positive correlations. Notably, the correlation strength of surpasses that of , indicating its superior relevance to brain activity. However, it is important to note that in some measures, the correlation coefficient of is higher than that of , which highlights the complexity of the task and underscores the critical role of bridging features. Furthermore, our results indicate that our generated visual latent features achieve activation levels comparable to the ground truth (GT), emphasizing the effectiveness of our comparative learning approach and demonstrating well-aligned fMRI feature and image spaces.
8 Conclusion
In this paper, we introduce the innovative task of Recon3DMind and its accompanying large-scale dataset, fMRI-3D, across various settings for the first time. Technologically, we develop a novel three-stage framework that involves multiple brain regions, including those associated with human 3D vision, specifically designed for this task. This approach establishes new benchmarks in the field and demonstrates the feasibility of the task. Initially, our model proficiently extracts features from fMRI frames. In the second stage, it employs a diffusion module to transition these fMRI features into the vision domain. Finally, in the third stage, the vision features are transformed into 3D models using an advanced 3D generation model. Our comprehensive experimental results and analyses affirm the model’s effectiveness in accurately extracting fMRI features and converting them into their corresponding 3D objects. We also conduct an in-depth analysis of our proposed fMRI-3D dataset and the features extracted by MinD-3D, further validating both the quality of fMRI-3D and the effectiveness of MinD-3D. This pioneering work not only opens up a new avenue in neuroimaging and 3D reconstruction but also paves the way for future research aimed at a deeper understanding and visualization of neural representations in 3D vision.